40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 12 23:50:10 2011, Suresh, Update, RF System, POY55_Demod board Hardware change completed 
    Reply  Wed May 18 07:18:15 2011, Suresh, Update, RF System, POY55_Demod board Hardware change completed 
Message ID: 4737     Entry time: Wed May 18 07:18:15 2011     In reply to: 4708
Author: Suresh 
Type: Update 
Category: RF System 
Subject: POY55_Demod board Hardware change completed 

The ~350 Hz noted in the elog below was traced to an RF modulation of the 11 MHz sideband.  This modulation was set up in the Marconi which is currently supplying the 11 MHz local oscillator signal to the RF source.  lt was used during the MC length study completed last week by Valera and Ryan.  The frequency measured was 322 Hz.

As we do not require this any longer, I have switched off this modulation.




The Demod board with S. No. 022 (being used earlier as REFL11) has been modified.  It now has SCLF-65 as its input LP filter on the PD input line and a PQW-2-90 power splitter.  The unit functioning okay (I and Q signals are 90 deg apart.

The loss of Q output was traced to a possible loose solder joint and we now have both the I and Q signals after resoldering all components in the vicinity of U7 (Ref Schematic of D990511)

There is a strong oscillation around 350Hz present on I and Q signals of both REFL55_Demod and POY55_Demod.  Don't know the source. 

We have run out of power splitters to continue with the Demod board modification. We do not currently have an AS11_Demod board.  All the others are in place and ready for the I<->Q phase angle measurement.

In summary we now have the following Demod boards in place:

[ REFL11, POY11, REFL55, AS55, POY55, POY22, POY110]_Demod



ELOG V3.1.3-